OpenStudy (anonymous):

What value is a discontinuity of x squared plus 2 x plus 3, all over x squared minus x minus 12?

OpenStudy (anonymous):

A x = -1 B x = -2 C x = -3 D x = -4

OpenStudy (anonymous):

?

OpenStudy (mathmate):

Hint: For a rational function, discontinuity occurs at locations where the denominator equals zero: \(\large f(x)=\frac{x^2+2x+3}{x^2-x-12}=\frac{x^2+2x+3}{(x-4)(x+3)}\) The denominator has been factorized to help you. All you need to do is to determine at which points the denominator becomes zero. (note: the answer choices include only one of the two values).